# Homeostatic Adaptation of Optimal Population Codes under Metabolic Stress

### v1

@@ Appendix E Relation between Approximate Homeostasis Constraint and Optimality
Proposition E.1 (Framework in [5] Implies Approx. Homeostasis).
@@
Appendix F Parameters and notation
We summarize how we configure the parameters in the simulation (see Tab.˜3) and list the notation (see Tab.˜4) in this paper.
@@ Appendix F Parameters and notation
mean of firing probability $\mu_{F}$
variance of firing probability $\sigma_{F}^{2}$
@@ Appendix G Biophysical simulation
The above simulation was performed on the distributed cluster computer based on Intel(R) Xeon(R) and does not require any GPU usage. We distributed the simulation of each cell states $(v_{rest},g_{leak},g_{leak})$ to a core with individual 1G memory. Each simulation that executes 10,000 run for estimating the mean and variance of firing probability takes approximately 40 minutes.
@@ H.1 Energy cost
where $\kappa$ represents the level of signal activity. In our simulations, $\kappa$ is varied over the range from 90 to 150. A detailed analysis of how different values of $\kappa$ influence the noise term $\eta_{\kappa}$ along the optimal adaptation path is provided in App.˜J.
@@ H.2 Tuning curve width
Following [22], we determine the $\hat{g}_{syn}$ that yields $\mu_{F}=0.1$ for each state $(v_{rest},g_{leak})$. Using these values of $\hat{g}_{syn}$, we define a tuning curve that maps stimulus to synaptic conductance:
@@
where $\phi:=0.02$ is used to compensate for the mean firing rate since $\hat{g}_{syn}$ is determined according to $\mu_{F}=0.1$. The stimulus $s$ is defined over the range from 90-90 to $90$ to simulate orientation input. The corresponding tuning curve from stimulus to mean firing rate is obtained by applying $TC$ through the empirically derived mapping from $g_{syn}$ to $\mu_{F}$ based on simulation data. In contrast to [22], we compute the full width at half maximum (FWHM) of the wrapped curve, as it may deviate from a Gaussian profile. To estimate the width along the optimal path, we fit the FWHM values using a second-degree polynomial function in $v_{rest}$ and $g_{leak}$ and take the estimated values along the optimal states.
@@ H.3 Spike train statistics
  • The variance of the firing probability $\sigma^{2}_{F}$ can be approximated by a parabola of the mean $\mu_{F}$ : $\sigma^{2}_{F}\approx\eta_{\kappa}\mu_{F}(1-\mu_{F})$.
@@
In other word, $K\sim\textit{Pois}(\lambda T)$, where $T$ is a given time window. Since our simulation input is a single input, we can estimate the mean and variance of the firing probability $F$ by the following empirical distribution:
@@
The last approximation holds when neurons operate at the low firing probability region, where we make $\mu_{F}=0.1$ in our simulation. The eq.˜78 also indicates a dispersed Poisson model. In eq.˜78, we use a parabolic function to model the variance of the firing probability. As shown in Fig.˜5, the fitted parabolas align closely with the simulation results, demonstrating the accuracy of the approximation.
Figure 5: Fitted parabolas for the mean and variance of firing probability. We fit a parabolic relationship between the mean and variance of firing probability across different cellular states $(v_{rest},g_{leak})$. For visualization purposes, only every second value of $v_{rest}$ and $g_{leak}$ is shown.
@@ H.4 Fitting of width, energy, and noise
We fit the estimated width, energy, and noise in 2, 1, and 8 degree polynomial functions in $v_{rest}$ and $g_{leak}$, respectively (see Fig.˜6). The relative errors of these fittings are less than $2\%$. We select a degree 1 for energy fitting to enhance the stability of the numerical optimization in eq.˜15, owing to the linearity of the constraint, which aligns with the straight energy contour. The degrees for the width and noise fitting are determined based on the maximum relative error.
Figure 6: Surface fitting to width, total energy, and noise. (a, b) We use 2 degree polynomial function in $v_{rest}$ and $g_{leak}$ to fit the estimated FWHM, with the relative error provided. (c, d) We use 1 degree polynomial function in $v_{rest}$ and $g_{leak}$ to fit the estimated total energy, with the relative error provided. (e, f) We use 8 degree polynomial function in $v_{rest}$ and $g_{leak}$ to fit the estimated total energy, with the relative error provided.
@@ Appendix I Change of energy contours under varying activity level
The change of energy contours under varying $\kappa$ is shown in Fig.˜7. Note that underlying noise is independent of $\kappa$ but the angle of constant-energy lines changes, becoming steeper with higher cell activity. This results in different curvatures of the optimal path shape.
@@
Here, we investigate how the activity level $\kappa$ influences the density (defined as the inverse of tuning curve width) and the behavior of the noise dispersion curve $\eta_{\kappa}(\epsilon)$ along the optimal adaptations. Specifically, we vary $\kappa$ from 90 to 150 and fit the proposed models to the corresponding density and noise data as described in eqs.˜16 and 18.
The fitted parameters are shown in Fig.˜8a-e. To characterize their dependence on $\kappa$, we fit an affine relation to $(a_{1},a_{2})$ as a function of $\kappa$ (see dotted lines in Fig.˜8). Due to the presence of outliers as $\kappa<100$, we restrict the fitting to data with $\kappa\geq 100$. The resulting fits (dotted lines in Fig.˜8a, b) are given by:
$a_{1}(\kappa)$ $=-8.249\cdot 10^{-11}\kappa+1.784\cdot 10^{-8},$ (79)
$a_{2}(\kappa)$ $=3.786\cdot 10^{-4}\kappa+0.6931.$ (80)
Additionally, we fit polynomial models of degree 2, 4, and 2 to $b_{1}$, $b_{2}$, and $\eta_{0}$, respectively (see dashed lines in Fig.˜8c–e). The fitted expressions are:
$b_{1}(\kappa)$ $=3.959\cdot 10^{1}\kappa^{2}-6.201\cdot 10^{3}\kappa+3.333\cdot 10^{5},$ (81)
$b_{2}(\kappa)$ $=1.300\kappa^{4}-5.948\cdot 10^{2}\kappa^{3}+9.986\cdot 10^{4}\kappa^{2}-7.225\cdot 10^{6}\kappa+2.052\cdot 10^{8},$ (82)
$\eta_{0}(\kappa)$ $=-2.488\cdot 10^{3}\kappa^{2}+4.202\cdot 10^{5}\kappa+2.052\cdot 10^{9}.$ (83)
The fitted trends begin to deviate noticeably for $\kappa>130$, indicating a potential cutoff beyond which the fitted $\eta_{\kappa}$ becomes less reliable. We further visualize the consequences of these deviations in the resulting density and noise dispersion curves. As shown in Fig.˜8f, the predicted density values increasingly deviate from simulation results as $\kappa$ decreases. In contrast, the noise dispersion curves (see Fig.˜8g) diverge more significantly at higher $\kappa$, with the deviation at $\kappa=150$ being more pronounced than at $\kappa=90$. It should note that the values of $\vec{c}(\kappa)$ can be easily derived from the values of $\vec{a}$ and $\vec{b}$.
Figure 8: Relationship between $\kappa$ and the fitting parameters in the energy budget and the noise dispersion model as in eqs.˜16 and 18. (a, b) To represent the energy budget in ATP/s units, we fit the parameters $(a_{1},a_{2})$ as an affine function of $\kappa$. (c–e) To model the noise in ATP/s units, we fit the parameters $(b_{1},b_{2},\eta_{0})$ with polynomials of $\kappa$. Dotted lines indicate polynomial fits of degree 2, 4, and 2 to $b_{1}$, $b_{2}$, and $\eta_{0}$, respectively. (f, g) The resulting energy budget and noise dispersion curves (solid lines) generated using the fitted expressions are compared with the simulation data. For visualization purposes, we display every fourth value of $\kappa$ in the plotted results.
@@
We compare the our analytical solution to [5, 30] under the different objectives (see. Tabs.˜5, 6 and 7).
@@
Ours
Ganguli &
Simoncelli [5]
Wang et al. [30]
Density
(tuning width)-1
$d(s)$
$E^{\frac{1}{\alpha}}R(s)^{-1}\,p(s)$ $Np(s)$ $Cp(s)$
Gain
$g(s)$
$E^{\frac{1}{\alpha}}$ $R$ 11
Fisher
information
$FI(s)$
$\frac{E^{\frac{3}{\alpha}}\,p(s)^{2}}{\eta_{\kappa}(E)\,R(s)^{2}}$ $N^{2}Rp(s)$ $C^{2}p(s)^{2}$
@@
Ours
Ganguli &
Simoncelli [5]
Wang et al. [30]
Density
(tuning width)-1
$d(s)$
$\propto E^{\frac{1}{\alpha}}R(s)^{\frac{-\alpha-1}{\alpha+3}}p(s)^{\frac{\alpha+1}{\alpha+3}}$ $\propto Np(s)^{\frac{1}{2}}$ $\propto Cp(s)^{\frac{1}{3}}$
Gain
$g(s)$
$\propto E^{\frac{1}{\alpha}}R(s)^{\frac{2}{\alpha+3}}p(s)^{\frac{-2}{\alpha+3}}$ $\propto Rp(s)^{\frac{-1}{2}}$ 11
Fisher
information
$FI(s)$
$\propto\frac{E^{\frac{3}{\alpha}}p(s)^{\frac{2\alpha}{\alpha+3}}}{\eta_{\kappa}(E)R(s)^{\frac{2\alpha}{\alpha+3}}}$ $\propto N^{2}Rp(s)^{\frac{1}{2}}$ $C^{2}p(s)^{\frac{2}{3}}$
@@
Ours
Ganguli &
Simoncelli [5]
Wang et al. [30]
Density
(tuning width)-1
$d(s)$
$\propto E^{\frac{1}{\alpha}}R(s)^{\frac{\alpha-\beta}{3\beta-\alpha}}p(s)^{\frac{\beta-\alpha}{3\beta-\alpha}}$ $\propto Np(s)^{\frac{\beta-1}{3\beta-1}}$ $\propto Cp(s)^{\frac{1}{1-2\beta}}$
Gain
$g(s)$
$\propto E^{\frac{1}{\alpha}}R(s)^{\frac{2\beta}{3\beta-\alpha}}p(s)^{\frac{-2\beta}{3\beta-\alpha}}$ $\propto Rp(s)^{\frac{2\beta}{1-3\beta}}$ 11
Fisher
information
$FI(s)$
$\propto\frac{E^{\frac{3}{\alpha}}p(s)^{\frac{-2\alpha}{3\beta-\alpha}}}{\eta_{\kappa}(E)R(s)^{\frac{2\alpha}{\alpha-3\beta}}}$ $\propto N^{2}Rp(s)^{\frac{2}{1-3\beta}}$ $C^{2}p(s)^{\frac{2}{1-2\beta}}$
@@
In Sec.˜5, we aim to find the parameters $\vec{a}$ for simultaneously accomplish 32% widening and 29% reduction in ATP consumption. Combining with the density analytical solution in our model, the 32% widening implies
@@
Given a fixed tuning curve width derived from the fixed population size, we reduce the firing rate according ($R$) in [5] according to the decrease in the maximum mean firing rate in the mouse neuron. Due to the gaussian assumption, 32% widening, and a uniform prior, we can calculate the decrease in the maximum mean firing rate by
@@
Similarly, due to 32% widening and a uniform prior, we can calculate the decrease in the coding capacity in [30] by
@@
Here, we analyze changes in firing rate under different coding objectives and prior distributions (see Fig.˜9). Across all coding objectives, we adopt a consistent configuration: $b/(a\epsilon_{ctr})=0.19625$, a 29% reduction in energy expenditure, and $\alpha=1$, consistent with the settings used in Sec.˜5. Additionally, we fix $E=6$ and $R(s)=1$ for all cases, as these parameters yield a control tuning width of 35 degrees under a uniform prior — following the simulation conditions reported in [22]. After generating the population codes according to Tab.˜1, we extract the tuning curve corresponding to the neuron with preferred stimulus $s_{n}=90^{\circ}$. Each extracted tuning curve is normalized by the peak value of its respective control condition, a transformation that does not affect the analysis of firing rate deviation.
@@
Fig.˜9a shows the uniform prior assumed in Sec.˜5, along with the infomax case (Fig.˜9b reiterates Fig.˜4b). Here we illustrate that, under constant $R(s)=R$, all objective functions lead to the same original tuning curves and adaptations, all with perfect homeostasis.
@@
In addition to the uniform prior, we also consider as a non-uniform prior the distribution of edge orientations in natural scenes, as reported in [8]. When replacing the uniform prior with the non-uniform one, the deviation in mean firing rate (compared to each control condition, shown in dotted purple) remains within 2% across infomax, discrimax, and $L_{1}$ error objectives. As shown in Fig.˜9g and h, the tuning curves deviate from a Gaussian profile due to the dependence of the gain function $g(s)$ on the prior distribution for $L_{p}$ norm losses (see Tabs.˜6 and 7).

### v2

@@ Appendix E Relation between Approximate Homeostasis Constraint and Optimality
Proposition E.1 (Framework in ganguli2010implicit Implies Approx. Homeostasis).
@@
Appendix F Simulation overview
Recall that current reversal is a significant energy cost, so that at a high level each of these parameters has the following effect:
Resting potential $v_{rest}$:
When $v_{rest}$ is less negative, receiving spikes is cheaper, because it takes less of a voltage change (and therefore less current) to reach the spiking threshold. Input current must be reversed by ion pumps using ATP. Perhaps counterintuitively, a less-negative $v_{rest}$ also makes sitting at rest more expensive. This is because a less negative $v_{rest}$ shifts the balance between the reversal potentials of Na+ and K+, which requires counteracting a higher influx of Na+ and costs more energy to maintain (see App. G for details and relevant equations). This cost trade-off between maintaining rest potential and reversing the current associated with incoming spikes is why the optimal paths shown in Fig. 3g vary in shape as a function of the input activity level of the cell.
Leak channel conductance $g_{leak}$:
Leak channels are a major factor in the overall membrane conductance. Using $V=IR=I/g$ we see that this mechanism also reduces the total current needed to make a voltage change in the cell.
Synaptic conductance $g_{syn}$:
This prevents the cell from firing excessively by counteracting the changes in $v_{rest}$ and $g_{leak}$ that reduce the energy cost of reaching spiking threshold. It scales down the impact of each incoming spike so that the mean firing rate of the cell is maintained. Essentially, for a cell that is twice as ready to fire, we adjust the synaptic conductance so that each incoming spike only creates half the change it usually would. This rescaled state is not identical in function to the original because the smaller distance to threshold, particularly in combination with the higher voltage variability also observed in this state, results in more accidental output spikes. (We will illustrate this effect with a figure of simulated voltage traces, showing a cell that requires three sequential input spikes to fire: in the well-fed case, two large voltage steps will not cause an output spike even in the presence of noise, but in the starving case, two smaller steps result in a smaller gap from threshold, which can be overcome by noise and cause an unintended output spike.) Note that this variable is how we inject a specific tuning curve into the model, following padamsey2022neocortex.
We additionally summarize how we configure the parameters in the simulation (see Tab. 3) and list the notation (see Tab. 4) in this paper.
@@ Appendix F Simulation overview
mean of spike count $\mu_{F}$
variance of spike count $\sigma_{F}^{2}$
@@ Appendix G Biophysical simulation
The overview of the simulation pipeline is shown in Fig. 2. We focus on three key cellular properties — resting potential $v_{rest}$, leak channel conductance $g_{leak}$, and synaptic conductance $g_{syn}$ — as the primary variables in our simulation, based on findings by padamsey2022neocortex, who reported that neurons adapt their encoding approach through changes in these parameters.
@@
The above simulation was performed on the distributed cluster computer based on Intel(R) Xeon(R) and does not require any GPU usage. We distributed the simulation of each cell states $(v_{rest},g_{leak},g_{leak})$ to a core with individual 1G memory. Each simulation that executes 10,000 run for estimating the mean and variance of spike count takes approximately 40 minutes.
@@ H.1 Energy cost
where $\kappa$ represents the level of signal activity. In our simulations, $\kappa$ is varied over the range from 90 to 150. A detailed analysis of how different values of $\kappa$ influence the noise term $\eta_{\kappa}$ along the optimal adaptation path is provided in App. J.
@@ H.2 Tuning curve width
Following padamsey2022neocortex, we determine the $\hat{g}_{syn}$ that yields $\mu_{F}=0.1$ for each state $(v_{rest},g_{leak})$. Using these values of $\hat{g}_{syn}$, we define a tuning curve that maps stimulus to synaptic conductance:
@@
where $\phi:=0.02$ is used to compensate for the mean firing rate since $\hat{g}_{syn}$ is determined according to $\mu_{F}=0.1$. The stimulus $s$ is defined over the range from 90-90 to $90$ to simulate orientation input. The corresponding tuning curve from stimulus to mean firing rate is obtained by applying $TC$ through the empirically derived mapping from $g_{syn}$ to $\mu_{F}$ based on simulation data. In contrast to padamsey2022neocortex, we compute the full width at half maximum (FWHM) of the wrapped curve, as it may deviate from a Gaussian profile. To estimate the width along the optimal path, we fit the FWHM values using a second-degree polynomial function in $v_{rest}$ and $g_{leak}$ and take the estimated values along the optimal states.
@@ H.3 Spike train statistics
  • The variance of the spike count $\sigma^{2}_{F}$ can be approximated by a parabola of the mean $\mu_{F}$ : $\sigma^{2}_{F}\approx\eta_{\kappa}\mu_{F}(1-\mu_{F})$.
@@
In other word, $K\sim\textit{Pois}(\lambda T)$, where $T$ is a given time window. Since our simulation input is a single input, we can estimate the mean and variance of spike count $F$ by the following empirical distribution:
@@
The last approximation holds when neurons operate at the low spiking region, where we make $\mu_{F}=0.1$ in our simulation. The eq. 77 also indicates a dispersed Poisson model. In eq. 77, we use a parabolic function to model the variance of spike count. As shown in Fig. 5, the fitted parabolas align closely with the simulation results, demonstrating the accuracy of the approximation.
Figure 5: Fitted parabolas for the mean and variance of spike count. We fit a parabolic relationship between the mean and variance of spike count across different cellular states $(v_{rest},g_{leak})$. For visualization purposes, only every second value of $v_{rest}$ and $g_{leak}$ is shown.
@@ H.4 Fitting of width, energy, and noise
We fit the estimated width and energy in 2 and 1 degree polynomial functions in $v_{rest}$ and $g_{leak}$, respectively (see Fig. 6). Additionally, the estimated noise is fitted with
$P_{4}(v_{rest},g_{leak})e^{w_{1}v_{rest}+w_{2}g_{leak}},$ (78)
where $P_{4}$ denotes the degree-4 polynomial in $v_{rest}$ and $g_{leak}$, and $w_{1}$ and $w_{2}$ are the learning weights. Note that we additionally impose a non-negative constraint on the noise fitting parameters, further mitigating potential overfitting. The relative errors of these fittings are less than $2\%$. We select a degree 1 for energy fitting to enhance the stability of the numerical optimization in eq. 15, owing to the linearity of the constraint, which aligns with the straight energy contour. The degrees for the width and noise fitting are determined based on the maximum relative error.
Figure 6: Surface fitting to width, total energy, and noise. (a, b) We use 2 degree polynomial function in $v_{rest}$ and $g_{leak}$ to fit the estimated FWHM, with the relative error provided. (c, d) We use 1 degree polynomial function in $v_{rest}$ and $g_{leak}$ to fit the estimated total energy, with the relative error provided. (e, f) We use eq. 78 to fit the estimated total energy, with the relative error provided.
@@ Appendix I Change of energy contours under varying activity level
The change of energy contours under varying $\kappa$ is shown in Fig. 7. Note that underlying noise is independent of $\kappa$ but the angle of constant-energy lines changes, becoming steeper with higher cell activity. This results in different curvatures of the optimal path shape.
@@
Here, we investigate how the activity level $\kappa$ influences the density (defined as the inverse of tuning curve width) and the behavior of the noise dispersion curve $\eta_{\kappa}(\epsilon)$ along the optimal adaptations. Specifically, we vary $\kappa$ from 90 to 150 and fit the proposed models to the corresponding density and noise data as described in Sec. 4.1.
The fitted parameters are shown in Fig. 8a-e. To characterize their dependence on $\kappa$, we fit an affine relation to $(a_{1},a_{2})$ as a function of $\kappa$ (see dotted lines in Fig. 8). Due to the presence of outliers as $\kappa<100$, we restrict the fitting to data with $\kappa\geq 100$. The resulting fits (dotted lines in Fig. 8a, b) are given by:
$a_{1}(\kappa)$ $=-6.193\cdot 10^{-11}\kappa+1.462\cdot 10^{-8},$ (79)
$a_{2}(\kappa)$ $=1.626\cdot 10^{-4}\kappa+0.7452.$ (80)
Additionally, we fit polynomial models of degree 2, 2, and 2 to $b_{1}$, $b_{2}$, and $\eta_{0}$, respectively (see dashed lines in Fig. 8c–e). The fitted expressions are:
$b_{1}(\kappa)$ $=3.093\cdot 10^{1}\kappa^{2}-3.922\cdot 10^{3}\kappa+2.013\cdot 10^{5},$ (81)
$b_{2}(\kappa)$ $=-5.996\cdot 10^{2}\kappa^{2}+1.888\cdot 10^{5}\kappa+3.370\cdot 10^{6},$ (82)
$\eta_{0}(\kappa)$ $=-6.102\cdot 10^{-7}\kappa^{2}+6.004\cdot 10^{-5}\kappa+9.934\cdot 10^{-1}.$ (83)
We further visualize the consequences of these deviations in the resulting density and noise dispersion curves. As shown in Fig. 8f, g, the predicted density and noise dispersion values increasingly deviate from simulation results as $\kappa$ decreases. It should note that the values of $\vec{c}(\kappa)$ can be easily derived from the values of $\vec{a}$ and $\vec{b}$.
Figure 8: Relationship between $\kappa$ and the fitting parameters in the energy budget and the noise dispersion model as in eqs. 16 and 18. (a, b) To represent the energy budget in ATP/s units, we fit the parameters $(a_{1},a_{2})$ as an affine function of $\kappa$. (c–e) To model the noise in ATP/s units, we fit the parameters $(b_{1},b_{2},\eta_{0})$ with polynomials of $\kappa$. Dotted lines indicate polynomial fits of degree 2, 2, and 2 to $b_{1}$, $b_{2}$, and $\eta_{0}$, respectively. (f, g) The resulting energy budget and noise dispersion curves (solid lines) generated using the fitted expressions are compared with the simulation data. For visualization purposes, we display every fourth value of $\kappa$ in the plotted results.
@@
We compare the our analytical solution to ganguli2010implicit; wang2016efficient under the different objectives (see. Tabs. 5, 6 and 7).
@@
Ours Ganguli & Simoncelli (2010) Wang et al. (2016a)
Density (tuning width)-1 $d(s)$ $E^{\tfrac{1}{\alpha}}R(s)^{-1}p(s)$ $Np(s)$ $Cp(s)$
Gain $g(s)$ $E^{\tfrac{1}{\alpha}}$ $R$ 11
Fisher information $FI(s)$ $\dfrac{E^{\tfrac{3}{\alpha}}\,p(s)^{2}}{\eta_{\kappa}(E)\,R(s)^{2}}$ $N^{2}Rp(s)$ $C^{2}p(s)^{2}$
@@
Ours Ganguli & Simoncelli (2010) Wang et al. (2016a)
Density (tuning width)-1 $d(s)$ $\propto E^{\tfrac{1}{\alpha}}R(s)^{\tfrac{-\alpha-1}{\alpha+3}}p(s)^{\tfrac{\alpha+1}{\alpha+3}}$ $\propto Np(s)^{\tfrac{1}{2}}$ $\propto Cp(s)^{\tfrac{1}{3}}$
Gain $g(s)$ $\propto E^{\tfrac{1}{\alpha}}R(s)^{\tfrac{2}{\alpha+3}}p(s)^{\tfrac{-2}{\alpha+3}}$ $\propto Rp(s)^{\tfrac{-1}{2}}$ 11
Fisher information $FI(s)$ $\propto\dfrac{E^{\tfrac{3}{\alpha}}p(s)^{\tfrac{2\alpha}{\alpha+3}}}{\eta_{\kappa}(E)R(s)^{\tfrac{2\alpha}{\alpha+3}}}$ $\propto N^{2}Rp(s)^{\tfrac{1}{2}}$ $C^{2}p(s)^{\tfrac{2}{3}}$
@@
Ours Ganguli & Simoncelli (2010) Wang et al. (2016a)
Density (tuning width)-1 $d(s)$ $\propto E^{\tfrac{1}{\alpha}}R(s)^{\tfrac{\alpha-\beta}{3\beta-\alpha}}p(s)^{\tfrac{\beta-\alpha}{3\beta-\alpha}}$ $\propto Np(s)^{\tfrac{\beta-1}{3\beta-1}}$ $\propto Cp(s)^{\tfrac{1}{1-2\beta}}$
Gain $g(s)$ $\propto E^{\tfrac{1}{\alpha}}R(s)^{\tfrac{2\beta}{3\beta-\alpha}}p(s)^{\tfrac{-2\beta}{3\beta-\alpha}}$ $\propto Rp(s)^{\tfrac{2\beta}{1-3\beta}}$ 11
Fisher information $FI(s)$ $\propto\dfrac{E^{\tfrac{3}{\alpha}}p(s)^{\tfrac{-2\alpha}{3\beta-\alpha}}}{\eta_{\kappa}(E)R(s)^{\tfrac{2\alpha}{\alpha-3\beta}}}$ $\propto N^{2}Rp(s)^{\tfrac{2}{1-3\beta}}$ $C^{2}p(s)^{\tfrac{2}{1-2\beta}}$
@@
In Sec. 5, we aim to find the parameters $\vec{a}$ for simultaneously accomplish 32% widening and 29% reduction in ATP consumption. Combining with the density analytical solution in our model, the 32% widening implies
@@
Given a fixed tuning curve width derived from the fixed population size, we reduce the firing rate according ($R$) in ganguli2010implicit according to the decrease in the maximum mean firing rate in the mouse neuron. Due to the gaussian assumption, 32% widening, and a uniform prior, we can calculate the decrease in the maximum mean firing rate by
@@
Similarly, due to 32% widening and a uniform prior, we can calculate the decrease in the coding capacity in wang2016efficient by
@@
Here, we analyze changes in firing rate under different coding objectives and prior distributions (see Fig. 9). Across all coding objectives, we adopt a consistent configuration: $b/(a\epsilon_{ctr})=0.19625$, a 29% reduction in energy expenditure, and $\alpha=1$, consistent with the settings used in Sec. 5. Additionally, we fix $E=6$ and $R(s)=1$ for all cases, as these parameters yield a control tuning width of 35 degrees under a uniform prior — following the simulation conditions reported in padamsey2022neocortex. After generating the population codes according to Tab. 1, we extract the tuning curve corresponding to the neuron with preferred stimulus $s_{n}=90^{\circ}$. Each extracted tuning curve is normalized by the peak value of its respective control condition, a transformation that does not affect the analysis of firing rate deviation.
@@
Fig. 9a shows the uniform prior assumed in Sec. 5, along with the infomax case (Fig. 9b reiterates Fig. 4b). Here we illustrate that, under constant $R(s)=R$, all objective functions lead to the same original tuning curves and adaptations, all with perfect homeostasis.
@@
In addition to the uniform prior, we also consider as a non-uniform prior the distribution of edge orientations in natural scenes, as reported in girshick2011cardinal. When replacing the uniform prior with the non-uniform one, the deviation in mean firing rate (compared to each control condition, shown in dotted purple) remains within .2% across infomax, discrimax, and $L_{1}$ error objectives. As shown in Fig. 9g and h, the tuning curves deviate from a Gaussian profile due to the dependence of the gain function $g(s)$ on the prior distribution for $L_{p}$ norm losses (see Tabs. 6 and 7).